40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 26 22:45:14 2013, manasa, Update, General, Vent progress 
    Reply  Tue Jul 30 00:21:15 2013, Koji, Update, General, Vent progress 
       Reply  Tue Jul 30 11:30:55 2013, manasa, Update, General, ITMX oplev 
          Reply  Tue Jul 30 13:41:35 2013, Koji, Update, General, ITMX oplev 
             Reply  Tue Jul 30 15:38:24 2013, manasa, Update, General, ITMX oplev 
Message ID: 8929     Entry time: Fri Jul 26 22:45:14 2013     Reply to this: 8936
Author: manasa 
Type: Update 
Category: General 
Subject: Vent progress 

[Koji, Manasa]

The arms were brought back to resonance after yesterday's vent work.
DCPD gains were changed (TRX gain from -0.002 to -0.04; TRY gain changed from 0.002 to 0.04) to lock the arm with low power. 
 
X arm and Y arm were locked for IR.

We then decided to align IPANG because the input pointing seemed good - the right time to fix IPANG. The IPANG spot at the last in-vac steering mirror was aligned to be an inch low in pitch.
 
We then aligned SRC. SR2 and SR3 were adjusted to center the beam at SRM. SRM was aligned to the retro-reflected beam until we saw flashes. But this position of SRM is not at the good OSEM range. We should correct the SRM suspensions. We postponed this activity for next week and went ahead to look at the status of the AS beam.
 
Looking through the window and using IR viewer were not convincing enough. We will have to get the access connector out on Monday to do the AS alignment.
We then closed the chamber with light doors and locked the arm cavities again. We brought SRC and PRC to resonance and saw strong flashes at the BS_PRM camera. 
 
To do:
SRM need to be moved in order to remove the big bias in yaw
SRM OSEMS need to be adjusted
Access connector should come out
AS needs to be aligned satisfactorily
ITMX oplev steering mirrors in vacuum should be checked.
ELOG V3.1.3-